First meeting RTGF (Red Temática de Geometría y Física )

Miraflores de la Sierra, Madrid 24–28 November 2003

List of participants

• Luis Álvarez-Cónsul (CSIC, Madrid), l.alvarez-consul@maths.bath.ac.uk

• Pablo Angulo Ardoy (UCM, Madrid), panguloa@mat.ucm.es

• Marta Aparicio Arroyo (Univ. Salamanca), fataliax@hotmail.com

• Jaume Amorós (UPC, Barcelona), jaume.amoros@upc.es

• Adil Belhaj (IFT, CSIC-UAM, Madrid), adil.belhaj@uam.es

• Luis J. Boya (Univ. Zaragoza), luisjo@unizar.es

• Volker Braungart (UPC, Barcelona), volker-braungart@upc.es

• José F. Cariñena (Univ. Zaragoza), jfc@unizar.es

• Laura Costa (Univ. Barcelona), costa@mat.ub.es

• Luis Fernández (Univ. de los Andes, Bogotá), lfernand@uniandes.edu.co

• Marisa Fernández (Univ. Pais Vasco, Bilbao), mtpferol@lg.ehu.es

• Oscar García-Prada (CSIC, Madrid), oscar.garcia@uam.es

• María P. Garcia del Moral Zabala (IFT, CSIC-UAM, Madrid), marigmz@hotmail.com • César Gómez (IFT, CSIC-UAM, Madrid), cesar.gomez@uam.es

• Tomás L. Gómez (UCM, Madrid), tgomez@mat.ucm.es

• Daniel Hernández Ruipérez (Univ. Salamanca), ruiperez@usal.es

• Nigel Hitchin (Univ. Oxford), hitchin@maths.ox.ac.uk

• Alberto Ibort (Univ. Carlos III, Madrid), albertoi@math.uc3m.es

• Manuel de León (CSIC, Madrid) mdeleon@imaff.cfmac.csic.es

• Marina Logares (UAM, Madrid), marina.logares@uam.es

• Juan J. Manjarin (IFT, CSIC-UAM, Madrid), juanjose.manjarin@uam.es

• David Martínez (Univ. Carlos III, Madrid), dmtorres@math.uc3m.es

• Ruben Minasian (École Polytéchnique, Palaiseau), ruben@cpht.polytechnique.fr • Ignasi Mundet i Riera (Univ. Barcelona), ignasi.mundet@uam.es

• Vicente Muñoz (UAM, Madrid): Formality of symplectic manifolds, vicente.munoz@uam.es • Peter Newstead (Univ. Liverpool and CSIC, Madrid), newstead@liv.ac.uk

• Fabio Pioli (CSIC, Madrid), fabio.p@imaff.cfmac.csic.es

• Christoph Schweigert (Univ. Hamburg), fmdv005@math.uni-hamburg.de

• Ignacio Sols (UCM, Madrid), sols@mat.ucm.es

• Carlos Tejero Prieto (Univ. Salamanca), carlost@usal.es

• Richard Thomas (Imperial College, London), richard.thomas@imperial.ac.uk

• Angel Uranga (IFT, CSIC-UAM, Madrid), angel.uranga@uam.es

• Guennadi Volkov (UAM, Madrid), volkov@delta.ft.uam.es